OpenStudy (anonymous):

find indefinite integral of (1+ sqrt(x)^14)

OpenStudy (shubhamsrg):

what do you mean by sqrt(x)^14 ? can you draw it out ?

OpenStudy (shubhamsrg):

the question i mean,,for clarification..

OpenStudy (zzr0ck3r):

sqrt(x)^14 = x^(14/2) = x^7?

OpenStudy (anonymous):

\[(1+\sqrt{x})^{14}\]

OpenStudy (anonymous):

I am supposed to find the indefinite integral

OpenStudy (shubhamsrg):

i see..then let, 1+sqrt(x) = t => dx/2sqrt(x) = dt =>dx= 2 (t-1) dt making that substitution, you'll find that you are reduced to a much simpler expression to integrate..

OpenStudy (anonymous):

@shubhamsrg I used that substitution but the answer I got still wasn't right. @kropot72 thanks for trying but i dont know how to get the integral from that

OpenStudy (shubhamsrg):

can you please show your work ?

OpenStudy (anonymous):

sure, doing it now

OpenStudy (anonymous):